A machine shop produces heavy duty high endurance 20-inch rods that are meant for use in a variety of military grade equipment. On occasion, the machine malfunctions and produces a groove or a chisel cut mark somewhere on the rod. If such defective rods can be cut so that there is at least 15 consecutive inches without a groove, then the rods can be salvaged for other purposes. If the location of the groove on a rod is described by a uniform distribution, what is the probability that a defective rod can be salvaged?
Answer #1

The defective rod can be salvaged if the groove lies on the rod between 0 and 5 inches OR 15 and 20 inches

P(X ≤ 5) = 5/(20-0) = 0.25

P(X ≥ 15) = 5/(20-0) = 0.25

The probability that a defective rod can be salvaged = P(X ≤ 5) + P(X ≥ 15) = 0.25+0.25

The probability that a defective rod can be salvaged = 0.50